ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

12. Probability and Statistics

1991 MS Classification: 62D05 Taljaard, Monica, University of Port Elizabeth Weighting Adjustments for Non-response: The Importance of forming Homogeneous Weighting Classes. The primary method of dealing with survey unit non-response is to increase the weights of respondents within certain sub-groups (called weighting classes) of the sample to compensate for the missing responses. Population or sample weighting adjustments for non-response are applied by weighting respondent data proportionally to the sample or population proportions in each weighting class, depending upon whether or not the population sizes in each weighting class are known. The ability of sample and population weighted estimators to reduce non-response bias depends on the formation of weighting classes in which the assumption of uniform response probabilities is tenable. This is illustrated by means of a practical example.